Suppose a country's real GDP is $18 trillion and that population is 300 million. 

Instructions: Enter your answers as whole numbers. 


a. What is this country's real GDP per capita? 

Suppose that during the next 10 years, real GDP grows by half and population triples in the country. 


b. After 10 years have passed, what will be this country's real GDP per capita? 
Answer #1

1) We have Real GDP of country = $18 trillion or 18,000,000,000,000

Population of country = 300 million = 300,000,000

Per capita real GDP of country =Real GDP of country /Population of country 18,000,000,000,000 / 300,000,000 = $60,000

 

2) After 10 years, Real GDP = $18 trillion * 1.5 times = $27 trillion or $27,000,000,000,000

Population = 300 million * 3 = 900 million or 900,000,000

Per capita real GDP = [Real GDP / Population] = [27,000,000,000,000 / 900,000,000] = $30,000
Answer #2

SOLUTION :


1 trillion = 10^6 million.


a.


Real GDP per capita 

= Real GDP / Population 

= 18*10^6 million $  /  300 million

= 60000 ($) (ANSWER)


b.


After 10 years :


New real GDP = 18*10^6 * 1.5 = 27*10^6 million $ 

New population = 3*300 = 900 million


So,


Real GDP per capita after 10 years

= New real GDP / New population

= 27*10^6 million $ / 900 million.

= 30000 ($) (ANSWER).

answered by: Tulsiram Garg